40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 1210     Entry time: Tue Feb 24 05:17:04 2015
Author: Zach 
Type: DailyProgress 
Category: M2 ISS 
Subject: First real M2 test 

Tonight I succeeded in using the M2 ISS readout board and the 3-mm diodes to do some real intensity stabilization using the SiFi test setup.

First, I built a foam box to use as a temporary enclosure for the PMC and diodes until we get our real box finished:

There are holes for the input and REFL beams, and the diodes are held with makeshift mounts that clamp down on the sockets. Clearly, these aren't as stiff or stable as what we're having built, but they do the job for now. There is a steering mirror before the 50/50 BS so that the position on each diode can be adjusted separately with ease.

I didn't want to open the Chachi ISS box can of worms yet, so I just built my own temporary breadboard circuit. I had done some preliminary SR560 locking, so I knew roughly what I wanted, and I measured the modulator -> PD transfer function again today and verified that it was flat well above 100 kHz. I made a 2-stage pole/zero-style circuit, with a double (removable) pole at 300 Hz and a zero at 10 kHz to bring phase back to -90° around the target UGF of 100 kHz. It looks like this:

I wanted to DC couple, so I came up with an idea to pick off the stable 5-V bias supply from the M2 board and sum it with the (negative) output of the in-loop PD in the first stage of the servo. I had some current-related issues with the summer at first, but these went away when I increased the input resistors a bit (n.b., to fix the gain I had to change some other components, and as a result the controller TF is actually slightly different than shown above, but not much).

Hooking it up, it locked right away with the expected UGF of near 100 kHz (not yet measured, but inferred from the transfer functions and spectra). Here is the stabilization result:

As you can see, the out-of-loop signal is stabilized to the shot noise level (which is \sqrt{2} higher than the bare shot noise for half the beam due to the well-understood correlated noise imprinted by the loop) from about 5 kHz down to just below 100 Hz. Below this, there is clearly some differential environmental noise between the PDs. I did some beam scanning to try and minimize with some success, but not much. I'm not sure what the coherence below 20 Hz indicates---the in-loop signal is suppressed to below the measurement noise level, while the OOL signal exhibits excess differential noise, so I don't see why there should be any coherence.

In any case, this is a nice verification that:

  • The M2 readout board works with real optical signals
  • The intensity feedback system for the SiFi experiment works
  • The 3-mm diodes (save for the one bad one---see CRYO:1207) behave nicely, at least for these relatively low powers
  • The PMC -> PD scheme shows promise for our future tests with nicer hardware
ELOG V3.1.3-